Circular No. 7381

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 2000I AND 2000J A. V. Filippenko and A. L. Coil, University of California at Berkeley,

report that uncalibrated CCD spectra (range 330–1000 nm), obtained on Mar. 15 UT with the Shane 3-m reflector at Lick Observatory, reveal that SN 2000I (cf. IAUC 7362) is of type II. Broad Hα emission and weak Hα absorption are visible. The approximate recession velocity is 6400 km/s, based on narrow emission lines in the host galaxy. SN 2000J (ibid.) is also of type II; the noisy spectrum, obtained under poor conditions, is consistent with a recession velocity of nearly 12 000 km/s.

SUPERNOVA 2000P IN NGC 4965 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that further spectra of SN 2000P, taken by P. Berlind on Mar. 14.4 and 15.4 UT, show that it is a type-IIn supernova. Spectra taken in good seeing show that the observed narrow (FWHM < 500 km/s) Hα and Hβ emission (cf. IAUC 7379) is intrinsic to the supernova. The combination of the narrow emission with broad absorption features of Hβ and He I makes this spectrum unusual. These spectra are posted on the

World Wide Web at http://cfa-www.harvard.edu/cfa/oir/Research/
supernova/RecentSN.html.

V4642 SAGITTARII = NOVA SAGITTARII 2000 N. N. Samus, Sternberg Astronomical Institute, Moscow, on behalf of

IAU Commission 27, informs us that the designation V4642 Sgr has now been given to this nova (cf. IAUC 7362, 7363, etc.).

Visual magnitude estimates by A. Pearce, Nedlands, W. Australia: Feb. 16.815 UT, 12.8; 17.826, 12.4; Mar. 2.845, 12.6; 4.809, 12.7; 6.874, 12.7; 7.865, 12.7.

V1494 AQUILAE Visual magnitude estimates: Jan. 15.265 UT, 8.2 (B. H. Granslo, Fjell-

hamar, Norway); 28.252, 8.7 (P. Schmeer, Bischmisheim, Germany); Feb. 11.260, 8.6 (A. Pereira, Cabo da Roca, Portugal); 22.185, 9.5 (K. Hornoch, Lelekovice, Czech Rep.); 28.205, 9.6 (Schmeer); Mar. 2.238, 8.5 (Pereira); 6.876, 9.4 (Pearce); 12.194, 9.8 (Schmeer); 15.194, 8.9 (Schmeer).

2000 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green